OpenStudy (anonymous):

Solve System Equations: x and y –x + 2y = -4 2x – 3y = 8

OpenStudy (anonymous):

Your first step is going to be to eliminate one of the variables. In this case it is easiest to choose x first. Multiply the first equation by 2.

OpenStudy (anonymous):

Example of that would be?

OpenStudy (anonymous):

If you had two equations x + y = 3 and 2x + 3y = 7. Your first step in solving it would be to multiply the first equation by 2 to get 2x + 2y = 6. There are still more steps, I'm just trying to walk you through them one by one

OpenStudy (anonymous):

Ok so then whats the next step?

OpenStudy (anonymous):

The next step will be to either add or subtract the two equations to eliminate one of the variables. In the example I gave, you would subtract the 2x +2y = 6 from the 2x + 3y = 7 to get y = 1. In yours, you want to add the two together: -2x + 4y = -8 + 2x - 3y = 8 ___________________ 0x + y = 0

OpenStudy (anonymous):

so x=0?

OpenStudy (anonymous):

y = 0. 0* any number is always 0 so we don't know the value for x yet. The next step will be to put 0 into one of the equations for y and solve for x.

OpenStudy (anonymous):

Ok can you do that so I can have a visual?

OpenStudy (anonymous):

Sure. Give me a second while I type it all out for you

OpenStudy (anonymous):

0x + y = 0 is the same as just y = 0 Putting 0 for y into one of the equations gives you: -x + 2(0) = -4 Simplify: -x + 0 = -4 -x = -4 Divide both sides by negative one to remove the -: x = 4

OpenStudy (anonymous):

Ok I get it! Thanks man appreciate it!

OpenStudy (anonymous):

Not a problem :)